### v1

# GR-MG: Leveraging Partially Annotated Data via Multi-Modal Goal Conditioned Policy

Abstract
The robotics community has consistently aimed to achieve generalizable robot manipulation with flexible natural language instructions. One of the primary challenges is that obtaining robot data fully annotated with both actions and texts is time-consuming and labor-intensive. However, partially annotated data, such as human activity videos without action labels and robot play data without language labels, is much easier to collect. Can we leverage these data to enhance the generalization capability of robots? In this paper, we propose GR-MG, a novel method which supports conditioning on both a language instruction and a goal image. During training, GR-MG samples goal images from trajectories and conditions on both the text and the goal image or solely on the image when text is unavailable. During inference, where only the text is provided, GR-MG generates the goal image via a diffusion-based image-editing model and condition on both the text and the generated image. This approach enables GR-MG to leverage large amounts of partially annotated data while still using language to flexibly specify tasks. To generate accurate goal images, we propose a novel progress-guided goal image generation model which injects task progress information into the generation process, significantly improving the fidelity and the performance. In simulation experiments, GR-MG improves the average number of tasks completed in a row of 5 from 3.35 to 4.04. In real-robot experiments, GR-MG is able to perform 47 different tasks and improves the success rate from 62.5% to 75.0% and 42.4% to 57.6% in simple and generalization settings, respectively. Code and checkpoints will be available at the project page: https://gr-mg.github.io/.
Figure 1: Overview. GR-MG consists of two modules: the progress-guided goal image generation model and the multi-modal goal-conditioned policy. The former generates a goal image based on the current observation, language instruction, and task progress. The latter predicts task progress and actions based on the goal image produced by the former. Data without action labels can be used to train the goal image generation model, while the multi-modal goal-conditioned policy can be trained on robot play data that lacks language annotations. Fully annotated data will be used to train both modules.
I INTRODUCTION
The robotics research community is striving to achieve generalizable robot manipulation using language-based instructions. Among various methods, imitation learning from human demonstrations is one of the most promising endeavors[1, 2, 3, 4, 5]. However, human demonstration data are scarce. And the process of collecting human demonstrations with action and language labels is time-consuming and labor-intensive. On the other hand, robot play data without language labeling is a scalable source of data. It requires no video clipping or language labeling and can be collected without human constant supervision [6]. These data can be automatically labeled with hindsight goal relabeling [7]: any frame in a video can be used as a goal image to condition the policy for predicting actions to evolve towards this frame from previous frames. Furthermore, there is a substantial collection of text-annotated human activity videos without action labels on the internet. These data contain valuable information about how the agent should move to change the environment according to the language description. Can we develop a policy to effectively leverage all the above partially annotated data?
Previous methods have ventured into this domain, but most were limited to using data that lacked either language labels or action labels [4, 7, 8, 9, 10]. Recent initiatives introduce diffusion models for generating goal images [11] or future videos [12, 13]. The generated image or video is then used as inputs for goal-conditioned policies or inverse dynamics models to predict actions, enabling the use of all the above-mentioned partially annotated data.
However, these approaches often overlook crucial information, such as task progress, during the goal generation phase. This omission can lead to inaccurate generated goals which significantly affects the subsequent action predictions. Furthermore, these policies rely solely on images or videos, making them brittle in the case where the generated goal is inaccurate.
To tackle these issues, we introduce GR-MG, a model designed to support multi-modal goals. It comprises two modules: a progress-guided goal image generation model and a multi-modal goal-conditioned policy. Datasets without action labels (e.g., text-annotated internet videos) can be used to train the goal image generation model. Datasets without language labels (e.g., robot play data) can be used to train the policy, with a null string as the language condition. Given that robot manipulation is a sequential decision-making process, where the task progress information can enhance prediction accuracy, we incorporate a novel task progress condition into our goal image generation model. This significantly improves the performance of both goal image generation and thus action prediction. During the training of the policy, we sample goal images from trajectories and condition the policy on both the goal image and the text instruction or solely on the goal image if text is unavailable. During inference, the policy utilizes both the language instruction and the goal image generated from the goal image generation model. Since GR-MG is conditioned on both the text and the goal image, the policy can still rely on the language condition to guide action prediction even if the generated goal image is inaccurate, substantially improving the robustness of the model.
We perform extensive experiments in the challenging CALVIN simulation benchmark [14] and a real-robot platform. In CALVIN, without using additional partially annotated data, GR-MG significantly outperforms all the comparing state-of-the-art methods in a zero-shot generalization setting, improving the success rate of completing 1 and 5 tasks in a row from 93.8% to 96.8% and 41.2% to 64.4%, respectively. When we add robot play data into training, GR-MG achieves average length of 3.11 with only 10% of provided fully annotated data, which is even better than the competitive GR-1 [4] baseline using all the data. In real-robot experiments, we evaluate GR-MG in both simple and 3 challenging generalization settings. In total, GR-MG can perform 47 tasks and improves the average success rate from 62.5% to 75.0% in the simple setting and from 42.4% to 57.6% in generalization settings. In summary, the contribution of this paper is threefold:
  • We propose a novel Generative Robot Policy with Multi-modal Goals (GR-MG), which is able to leverage both text-annotated videos and robot play data during training.
  • We introduce the task progress condition in goal image generation, which substantially improves the accuracy of the generated goal image.
  • We perform extensive experiments and ablation studies in simulation and the real world to verify the effectiveness of GR-MG in both simple and challenging generalization settings.
Figure 2: Network Architecture. We use a diffusion model to generate the goal image based on the current image, language instruction, and task progress. The generated goal image is then sent to the multi-modal goal-conditioned policy, which takes as inputs the language instruction, sequences of visual observations, and robot states. The policy is a GPT-style transformer. We insert query tokens [PROG], [OBS], and [ACT] after the input tokens to predict actions, future images, and the progress of the current state, respectively. The progress is fed back to the goal generation model. In this figure, we show the setup with two camera views: a static view and a view captured from a camera mounted on the end-effector.
II Related Work
II-A Leveraging Multi-Modal Goals in Policy Learning
Language is probably the most flexible and intuitive way for a human to specify tasks for a robot [1, 2, 4, 5, 15]. However, language and visual signals exist in different domains. And the information contained in a language instruction may be too abstract for a visual policy to comprehend. A line of research explores goal-image conditioned policies [16]. But obtaining goal images during rollouts can be challenging. Recently, some studies have proposed leveraging both language and goal images as conditions during training. They utilize data with language labels alongside robot play data that without language labels [7, 17, 18, 3]. A common approach is to project both the goal image and the language into a shared latent space to condition the policy. However, only the language is used as the condition during inference. GR-MG differs from these methods in that it uses both languages and goal images during inference and training. The goal image is generated via a diffusion model from the language and current observation. The most related work to GR-MG is Susie [11]. It also uses a diffusion model to generate the goal image. The goal image is then fed into a diffusion policy [19] for action prediction. GR-MG follows a similar paradigm but differs in two key aspects. First, it introduces a novel task progress condition in goal image generation, significantly enhancing the accuracy of the generated goal image. Second, GR-MG conditions the policy using both the language and the goal image, whereas Susie relies solely on the generated goal images. This design enables GR-MG to address situations where the generated goal image is inaccurate, thereby improving performance in challenging out-of-distribution environments.
II-B Leveraging Partially Annotated Data in Policy Learning
Given that collecting fully annotated robot data with action and language labels is time-consuming, many studies turn to alternative sources of data with partial annotations. One approach involves learning useful representations from text-annotated internet videos, which are subsequently used for downstream policy learning [20, 21]. Recently, RT-2 [2] introduced a vision-language-action model that incorporates both robot data and large-scale vision-language data during training, demonstrating powerful generalization capabilities. Another popular method is to learn a video model or world model from internet video data to predict future videos [22, 13, 12, 8]. An inverse dynamics model or goal-conditioned policy is then employed to predict actions based on the predicted video. Recently, Wu et al. [4] proposed GR-1, which is pre-trained on large-scale text-annotated videos from the internet and then fine-tuned on robot trajectories with language labels. GR-MG differs from GR-1 in two aspects. First, GR-MG leverages robot play data during training, while GR-1 does not. Second, GR-MG utilizes both the language and the goal image as conditions, whereas GR-1 relies solely on the language. These differences significantly enhance GR-MG’s generalization capabilities, as we will demonstrate in Sec. IV.
III Method
We aim to learn a policy $\pi$ to address the problem of language-conditioned visual robot manipulation. Specifically, the policy takes as inputs the following. 1) A language instruction $l$ that describes the task to be accomplished. 2) Observation images $\mathbf{o}_{t-h:t}$ which include RGB images from the current and previous $h$ timesteps. A static global view camera and a hand-mounted camera are used to capture these images. 3) Robot states $\mathbf{s}_{t-h:t}$ which include the 6-Dof pose of the end-effector and the binary gripper state of the current and previous $h$ steps. The policy outputs an action trajectory $\mathbf{a}_{t}$ in an end-to-end manner:
$\mathbf{a}_{t}=\pi(l,\mathbf{o}_{t-h:t},\mathbf{s}_{t-h:t})$ (1)
III-A Data
The training of a language-conditioned policy requires fully annotated robot trajectories $\tau$ which contain both action/state labels and language labels:
$\tau=\{l,(\mathbf{o}_{1},\mathbf{s}_{1},\mathbf{a}_{1}),(\mathbf{o}_{2},\mathbf{s}_{2},\mathbf{a}_{2}),\ldots,(\mathbf{o}_{T},\mathbf{s}_{T},\mathbf{a}_{T})\}$ (2)
However, collecting fully annotated data is time-consuming. The motivation behind GR-MG is to utilize partially annotated data. One such data includes videos without action labels, which can be collected at scale from the internet and various public datasets [23, 24, 25]. Another type of partially annotated data consists of robot play data without language labels. These can be gathered on a large scale by allowing robots to execute a policy without constant human supervision [6].
III-B Network Architecture
The network architecture of GR-MG is illustrated in Fig. 2. It consists of two modules: a progress-guided goal image generation model and a multi-modal goal conditioned policy.
III-B1 Progress-guided Goal Image Generation Model
This module generates the goal image based on the current observation image, a language description of the task, and the task progress. Specifically, we use InstructPix2Pix [26], a diffusion-based image-editing model, as the network. Following the approach in Susie [11], we generate a sub-goal representing an intermediate state rather than the final state, updating this sub-goal at fixed time intervals during the rollout. However, unlike Susie which generates the goal image based solely on the current image and the language instruction, we incorporate task progress information in the image generation process. The insight is that robot manipulation is a sequential decision process instead of static image editing. Additionally, a current image can be ambiguous in the case where there are identical observation throughout the rollout of a task, e.g., moving an object back-and-forth. The progress information can provide valuable global information for the goal image generation model. During training, the progress of a frame can be easily obtained from its timestep in the trajectory (or video). During inference, such information is not accessible. To address this, we train the policy to predict task progress alongside the action trajectory, as we will outline below. To incorporate progress information into the generation model, we append a sentence to the task description, e.g., "pick up the red block. And 60% of the instruction has been completed.". We use T5-Base [27] to encode the text.
III-B2 Multi-modal Goal Conditioned Policy
Following GR-1 [4], we employ a similar GPT-style transformer network [28]. For completeness, we briefly review GR-1 here. GR-1 takes as input a language instruction, a sequence of observation images, and a sequence of robot states. The model then outputs an action and a prediction of the future image. The language is encoded using a text encoder [29]; images are encoded via MAE [30]; robot states are encoded through linear layers. To predict actions, GR-1 includes a [ACT] query token, and for future image prediction, it learns multiple [OBS] query tokens. The embeddings of different signals and the query tokens are sequenced by timestep and input into the transformer. For more details, we refer readers to [4].
The difference between GR-MG and GR-1 is threefold. Firstly, GR-MG is a multi-modal goal-conditioned policy which takes both a goal image and a language instruction as the condition. GR-1 only takes the language. To inject the goal image condition, we use MAE [30] to tokenize the goal image generated by the goal image generation model and append the tokens at the front of all the input tokens (Fig. 2). The goal image tokens can be attended by all the subsequent tokens for conditioning. Secondly, in order to predict task progress, a [PROG] query token is included as shown in Fig. 2. The output embedding of the token is passed through linear layers to regress the progress value. Finally, we follow recent work [31] to predict action trajectories instead of a single action as in GR-1 via a conditional variational autoencoder (cVAE) [32, 33]. Specifically, we use a VAE encoder to encode the action trajectory into a style vector embedding. We concatenate the style vector embedding, the output embedding of the [ACT] token and $k$ learnable token embeddings together and input them to a transformer to predict an action trajectory of $k$ steps.
III-C Training
For the training of goal image generation model, we follow the training setting in InstructPix2Pix [26] and train a noise prediction model as in DDPM [34]. We sample the frame which is $N$ steps away from the current frame in the trajectory/video as the goal image. As a rough task progress can already be very informative and can bring about more stable training, we discretize the task progress predicted from the policy from 0% to 100% into 10 bins before passing it to the goal image generation model.
For the training of multi-modal goal-conditioned policy, we initialize its weight with the pre-trained model weight derived from the generative video pre-training on Ego4d [25] in alignment with GR-1 [4]. Subsequently, leveraging our multi-modal goals, the policy can be first trained on data without language annotations, e.g., robot play data. Specifically, we input a null string for the text input in this stage. After that, we can further finetune the policy with fully annotated robot trajectories.
The input task progress in the training of the goal image generation model is computed from the timestep of the image in the video. The goal image in the training of the policy is sampled from the ground truth goal image from the trajectory. Therefore, the training of the two modules are independent, although they are interconnected during inference. We leave the investigation of co-training the two modules as future work.
III-D Inference
During inference, the task progress is initially set as zero. And the goal image generation model uses the current observation image, language instruction, and the task progress to generate goal images. The goal image is passed to the multi-modal goal-conditioned policy together with the language instruction, sequence of observation images, and sequence of robot states to predict action trajectories and task progress. As the goal image is set as $N$ steps away from the current image in training, we run the goal image generation model only every $n<N$ steps in a closed-loop manner for efficiency.
IV Experiments
We perform extensive experiments in a simulation benchmark and the real world to evaluate the performance of GR-MG. Throughout the experiments, we aim to answer three questions. 1) Can GR-MG perform multi-task learning and deal with challenging settings including unseen distractors, unseen backgrounds, and unseen objects? 2) Does incorporating robot play data enhance the performance of GR-MG, especially when fully annotated data are scarce? 3) How does the inclusion of videos without action labels improves the performance of GR-MG?
TABLE I: Results On Calvin ABC->D
Method Fully Annotated Data Use Play Data No. Instructions in a Row (1000 chains)
1 2 3 4 5 Avg. Len.
Hulc[18] 100% 41.8% 16.5% 5.7% 1.9% 1.1% 0.67±0.10
MDT[17] 100% 61.7% 40.6% 23.8% 14.7% 8.7% 1.54±0.04
Spil[35] 100% 74.2% 46.3% 27.6% 14.7% 8.0% 1.71
Roboflamingo[5] 100% $\times$ 82.4% 61.9% 46.6% 33.1% 23.5% 2.47
Susie[11] 100% 87.0% 69.0% 49.0% 38.0% 26.0% 2.69
GR-1[4] 100% $\times$ 85.4% 71.2% 59.6% 49.7% 40.1% 3.06
3D Diff Actor[36] 100% $\times$ 93.8% 80.3% 66.2% 53.3% 41.2% 3.35±0.04
GR-MG w/o image 100% $\times$ 91.0% 79.1% 67.8% 56.9% 47.7% 3.42±0.28
GR-MG w/o lang 100% $\times$ 91.8% 79.8% 68.9% 58.1% 48.1% 3.46±0.04
GR-MG w/o prog 100% $\times$ 94.1% 85.0% 75.2% 65.4% 56.3% 3.76±0.11
GR-MG (Ours) 100% $\times$ 96.8% 89.3% 81.5% 72.7% 64.4% 4.04±0.03
GR-1[4] 10% $\times$ 67.2% 37.1% 19.8% 10.8% 6.9% 1.41±0.06
GR-MG w/o play data 10% $\times$ 82.4% 60.8% 42.2% 28.7% 19.7% 2.33±0.04
GR-MG (Ours) 10% 90.3% 74.5% 61.2% 47.4% 37.5% 3.11±0.08
  • "100%" and "10%" denote the proportion of fully annotated data used for training. "w/o prog" indicates that task progress is not provided to the goal image generation model. "w/o lang" indicates that the policy uses only the goal image as the condition, with the language replaced by an empty string. "w/o image" indicates that the policy uses only the language as the condition, with the goal image replaced by a zero image.
Figure 3: Experiments. CALVIN benchmark consists of 34 tasks across four different environments. Real-robot experiments encompass 47 tasks, including pick-and-place operations, manipulation of articulated objects and pouring task.
IV-A CALVIN Benchmark Experiments
IV-A1 Experiment Settings
We first perform experiments in the challenging CALVIN [14] benchmark environment. CALVIN focuses on language-conditioned visual robot manipulation. It contains 34 tasks and four different environments, i.e., Env A, B, C, and D (Fig. 3). Each environment contains a Franka Emika Panda robot with a parallel-jaw gripper and a desk for performing different manipulation tasks. Real-robot rollout videos can be found in the supplementary material and the project page.
IV-A2 Generalization to Unseen Environments
[…]
We compare GR-MG with various state-of-the-art methods, including one that utilizes 3D signals [36]. The results are presented in Tab. I. Baseline results are taken from the original papers or provided by the authors. GR-MG outperforms all the comparing methods, establishing a new state-of-the-art. The average length reported in the last column of Tab. I serves as a comprehensive metric, indicating the average number of tasks the robot is able to accomplish in a row of five across the evaluated 1,000 sequences. GR-MG improves the average length from 3.35 to 4.04. Additionally, it increases the success rate of completing one and five tasks in a row from 93.8% to 96.8% and from 41.2% to 64.4%, respectively. These results demonstrate GR-MG’s strong multi-task learning capabilities and its ability to generalize to unseen environments.
Figure 4: Visualization of the generated goal images in CALVIN Benchmark. The generated images of GR-MG closely align with the ground truth. However, without the progress information during the generation process, the generated goal images can significantly diverge from the language instruction.
IV-A3 Data Scarcity
[…]
IV-A4 Ablation Studies
[…]
IV-B Real-Robot Experiments
Figure 5: Real-Robot Experiment Settings. We evaluate four settings in the real-robot experiments. 1) Simple: This setting aligns with the training dataset. 2) Unseen Distractor: Unseen distractors are added to the scene. 3) Unseen Background: The background is altered by placing two new tablecloths on the table. 4) Unseen Object: The manipulated objects are replaced with randomly chosen unseen objects. The language instructions are modified correspondingly.
Figure 6: Success Rates of Real-Robot Experiments. On the last column, we show the average success rate of the three unseen generalization settings.
IV-B1 Experiment Settings
To evaluate the performance of GR-MG in the real world, we perform experiments on a real robot platform. The platform consists of a Kinova Gen-3 robot arm equipped with a Robotiq 2F-85 parallel-jaw gripper and two cameras, i.e., one static camera for capturing the workspace and another camera mounted on the end-effector. We collected about 15k human demonstrations across 29 tasks, which include pick-and-place manipulation, articulated object manipulation, and pouring. See Fig. 3 for some examples. We use these data to train GR-MG.
In this experiment, we aim to verify the effectiveness of leveraging partially annotated data without action labels in the training of goal image generation model. Some robot datasets have action labels. But their embodiments and coordinate settings are different from the target platform, making them hard for direct usage in model training. However, if the language labels are provided, they can be used as partially annotated data without action labels. In this experiment, we train the goal image generation model with our real robot data, Something-Something-V2 [23], and the videos extracted from RT-1 dataset [1, 37]. We note that, in general, any video with text annotations can be included in the training.
We design four different settings to evaluate the model performance (see Fig. 5): Simple, Unseen Distractor, Unseen Background, and Unseen Object. In Simple, the scene is set to be similar to those in the training data. In Unseen Distractor, unseen distractors are added to the scene. This setting is challenging because the model may be distracted by these new elements, especially if they have similar colors. In Unseen Background, the background is modified by introducing two new tablecloths that were not present in the training data. In Unseen Object, the robot is instructed to manipulate objects that were not included in the training dataset. And the language instructions are adjusted accordingly, i.e., the language instructions are also unseen. In total, we evaluate 47 different language instructions: 29 of which were seen during training, while the remaining tasks were unseen. We compare the performance of GR-MG with GR-1 [4] and a variant of GR-MG, GR-MG w/o part. ann. data, which uses only the fully annotated data for training the goal image generation model. To ensure a fair comparison, given that GR-1 can also utilize text-annotated videos via generative video pre-training, we incorporate Something-Something-V2 and RT-1 into its pre-training.
IV-B2 Results
Results are shown in Fig. 6. Real-robot rollout videos can be found in the supplementary material and the project page. GR-MG outperforms GR-1 and GR-MG w/o part. ann. data in both simple and all the generalization settings. On average, GR-MG improves the average success rate of GR-1 from 62.5% to 75.0% in Simple setting and 42.4% to 57.6% in the three generalization settings. By comparing with GR-MG w/o part. ann. data, we can conclude that additional text-video data is crucial for enhancing the generalization capability of GR-MG. The failure case of GR-1 generally includes 1) picking wrong objects and 2) closing the gripper too early before arriving at the object. For GR-MG, the failure case mainly results from inaccurate actions, e.g., not able to move into the handle to open the oven in the oven opening task. We believe one reason is that GR-MG does not incorporate any depth information in the model. We plan to investigate incorporating depth information in future work.
To explore how partially annotated text-video data benefits GR-MG, we qualitatively compare the generated goal images between GR-MG and GR-MG w/o part. ann. data (Fig. 7). GR-MG produces high-quality and accurate goal images, while the generated goal images of GR-MG w/o part. ann. data is of comparatively low quality and do not adhere to the language instructions. Specifically, in Simple (Fig. 7 (a)), Unseen Distractor (Fig. 7 (b)), and Unseen Background (Fig. 7 (c)) settings, GR-MG w/o part. ann. data is more likely to generate images that fail to follow language instructions. In the Unseen Object setting (Fig. 7 (d)), the absence of additional partially annotated data increases the likelihood of hallucination: objects from the training set will unexpectedly appear in the generated goal images, even if they are not present in the current scene. These finding indicates that incorporating more text-annotated videos into training enhances 1) the understanding of the semantic in languages and 2) the robustness against out-of-distribution data.
Figure 7: Visualization of Generated Goal Images in Real-Robot Experiments. By utilizing additional text-annotated videos to train the goal image generation model, the generated images exhibit improved visual quality and better align with the language instructions.
V Conclusions
In this paper, we introduce GR-MG, a novel method that utilizes multi-modal goals to predict actions. GR-MG uses both a language and a goal image to condition the action prediction. It requires only the language as input and generate the goal image based on the language via a goal image generation model. This design enables GR-MG to effectively leverage large amounts of partially annotated data that are missing either language or action labels. To enhance the robustness of the generated goal images, we incorporate task progress information in the goal image generation model. GR-MG demonstrates exceptional performance in both simulation and real-world experiments, showing strong generalization across various out-of-distribution settings. Given the capability to utilize partially annotated data, we plan to scale up the training of both the goal image generation model and the policy in the future. Additionally, we plan to investigate integrating depth information to further improve the accuracy of action prediction.